(y+xz)p-(x+yz)q=x²-y²
1
Expert's answer
2020-04-20T11:58:35-0400


Given, "(y+xz)p-(x+yz)q = x^{2}-y^{2}".

This equation is of the form "Pp+Qq=R" (Lagrange's linear partial differential equation).

Here, "P = y+xz, Q=-(x+yz), R = x^{2}-y^{2}".


The subsidiary equations are

"\dfrac{dx}{P}=\dfrac{dy}{Q}=\dfrac{dz}{R}\\
\dfrac{dx}{y+xz}=\dfrac{dy}{-(x+yz)}=\dfrac{dz}{x^{2}-y^{2}}~~~~~~~~~~~-(1)".

Each ratio of (1) is equal to "\dfrac{xdx+ydy}{(x^{2}-y^{2})z} = \dfrac{dx+dy}{(1-z)(y-x)}".

Let us consider,

"\dfrac{xdx+ydy}{(x^{2}-y^{2})z} = \dfrac{dz}{x^{2}-y^{2}}\\
xdx+ydy=zdz\\
\text{Integrating, }\\ \int xdx + \int ydy = \int zdz\\
\dfrac{x^{2}}{2}+\dfrac{y^{2}}{2} =\dfrac{z^{2}}{2}+\dfrac{c_{1}}{2}\\
x^{2}+y^{2}-z^{2}=c_{1}"


Consider,

"\dfrac{dx+dy}{(1-z)(y-x)}=\dfrac{dz}{(x+y)(x-y)}\\
\dfrac{dx+dy}{1-z}=\dfrac{dz}{-(x+y)}\\
-(x+y)d(x+y)=(1-z)dz\\
\text{Integrating,}\\
-\int (x+y)d(x+y)=\int (1+z)dz\\
-\dfrac{(x+y)^2}{2}=\dfrac{(1-z)^{2}}{-2}+\dfrac{c_{2}}{2}\\
(1-z)^{2}-(x+y)^{2}=c_{2}"


The general solution is,

"\phi(c_{1},c_{2})=0\\
\phi\left(x^{2}+y^{2}-z^{2}, (1-z)^{2}-(x+y)^{2}\right)=0."